OpenStudy (anonymous):

Find the perimeter? There's a right triangle, with three sides, only two of the sides are given. I have to find X. The given sides are 12 and 16 I know the formula for finding perimeter but my answer doesn't match up with the multiple choice options. Help?

OpenStudy (wolf1728):

It's hard to answer that because we are not told if either of those sides is a hypotenuse. If it is both sides that = 12 & 16 then the hypotenuse² = 144 + 256 hypotenuse² = 400 hypotenuse = 20

OpenStudy (anonymous):

Okay, 20 was one of the options but I'll draw the triangle for you just in case. Can u explain to me how you found that answer? I get really confused 

OpenStudy (wolf1728):

Sure, I used the Pythagorean Theorem which states that the sum of the squares of the 2 sides of a right triangle equals the square of the hypotenuse. So 12² = 144 and 16² = 256 and so the hypotenuse² = 144 + 256 = 400 hypotenuse = square root of 400 which equals 20.

OpenStudy (anonymous):

Oh ok! See I was doing it completely differently. Thank you! I get what I was doing wrong, thanks so much!

OpenStudy (wolf1728):

Glad to help you out.